jhonyy9 (jhonyy9):

- let p and k prims,p>=2 and k>=2, - let a and b natural numbers ,a>=1 and b>=1, prove that for every n>=2 exist one a=(p-1)/2 and b=(k-1)/2 such that this equation n=a+b+1 is true . -for example : 2=(2-1)/2 +(2-1)/2 +1 2= 1/2 +1/2 +1 2=1+1 2=2 or 3=(3-1)/2 +(3-1)/2 +1 3=1+1+1 3=3 or 4=(5-1)/2 +(3-1)/2 +1 4=2+1+1 4=4

OpenStudy (anonymous):

n = (p-1)/2 + (k-1)/2 +1 = (p+k)/2. Thus, what the question is actually asking is to prove that are always an infinite number of primes>2 such that the sum of any two of them is even.But this is obviously true as any prime>2 is odd and the sum of two odd numbers is always even.